# A New Approach to Backtracking Counterfactual Explanations: A Causal Framework for Efficient Model Interpretability

### v1

@@ Abstract
Counterfactual explanations enhance interpretability by identifying alternative inputs that produce different outputs, offering localized insights into model decisions. However, traditional methods often neglect causal relationships, leading to unrealistic examples. While newer approaches integrate causality, they are computationally expensive. To address these challenges, we propose an efficient method based on backtracking counterfactuals that incorporates causal reasoning to generate actionable explanations. We first examine the limitations of existing methods and then introduce our novel approach and its features. We also explore the relationship between our method and previous techniques, demonstrating that it generalizes them in specific scenarios. Finally, experiments show that our method provides deeper insights into model outputs.
@@ 1 Introduction
Machine learning has become a core technology in areas such as healthcare, finance, and autonomous systems (Bhoi et al., 2024; Xie et al., 2024; Sancaktar et al., 2022). Although ML models are generally very effective, their limited interpretability is still a significant obstacle (Jethani et al., 2021). Understanding why a model generates a specific prediction is crucial for trust, fairness, and accountability (Miller, 2019; Zhang & Bareinboim, 2018; Von Kügelgen et al., 2022; Karimi et al., 2023). This need is especially clear in high-stakes domains like medical diagnosis or loan approval, where decisions can lead to serious consequences (Doshi-Velez & Kim, 2017).
@@
Despite their benefits, traditional counterfactual methods often overlook causal relationships between features, which can lead to impractical or unrealistic suggestions (Slack et al., 2021). For example, advising someone to lower their income while increasing savings ignores the causal dependency between these factors. This limitation reduces the practical value of such explanations. Causal algorithmic recourse (Karimi et al., 2021) incorporates Interventional Counterfactuals to produce more realistic outputs, but this approach is typically computationally expensive and difficult to scale.
Backtracking counterfactuals (Von Kügelgen et al., 2023) present a new way to define counterfactuals in causal inference. We propose a new framework for generating counterfactual explanations using backtracking counterfactuals. Our method combines causal reasoning with computational efficiency, enabling it to produce realistic and actionable explanations at scale. The key contributions of our work are as follows:
@@
  • We introduce a novel framework that leverages backtracking counterfactuals to provide actionable and meaningful explanations.
@@ 2.1 Structural Causal Models (SCMs)
where $\mathbf{X}_{\text{pa}(i)}$ are the parent variables (direct causes) of $X_{i}$, and $U_{i}$ are independent noise terms sampled from a distribution $P_{\mathbf{U}}$. These relationships are represented by a directed acyclic graph (DAG) $G$, which governs the observational distribution $P_{\mathbf{X}}^{\mathcal{C}}$.
@@
commonly referred to as the reduced-form expression. The function $\mathbf{F}$ translates the distribution of latent variables $\mathbf{U}$ into the distribution of observed variables $\mathbf{X}$. We assume causal sufficiency, implying no hidden confounders are present.
Additionally, we adopt a Bijective Generation Mechanism (Nasr-Esfahany et al., 2023), which assumes that $f_{i}(\mathbf{x}_{\text{pa}(i)},\cdot)$ is invertible for fixed $\mathbf{x}_{\text{pa}(i)}$. This ensures the existence of the inverse mapping $\mathbf{F}^{-1}$, allowing us to recover:
@@ 2.2 Interventional and Backtracking Counterfactuals
Both interventional and backtracking perspectives provide valuable insights into counterfactual reasoning but rely on distinct causal reasoning paradigms. Here, we only give a brief overview of these two approaches. Their precise definitions appear in Appendix A.
@@ 2.3 Problem Definition
The input $\mathbf{x}$ is assumed to follow a SCM $\mathcal{C}=(\mathbf{S},P_{\mathbf{U}})$, where the structural equations $\mathbf{S}$ are fully known. Formally, if $\mathbf{U}$ denotes the latent (noise) variables of the SCM, the input $\mathbf{X}$ is generated as $\mathbf{X}=\mathbf{F}(\mathbf{U})$, with the function $\mathbf{F}(.)$ explicitly defined. The components of $\mathbf{U}$ are mutually independent, and $\mathbf{F}(.)$ is invertible. The goal is to find a counterfactual input $\mathbf{x}^{\mathrm{CF}}$ that satisfies:
@@ 4 Our method
As discussed earlier, one of the main limitations of backtracking counterfactuals is their reliance on the conditional distribution $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$. This dependency arises because the choice of $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$ significantly influences the resulting counterfactuals, and its specification is left entirely to the algorithm. Such a distribution is essential when a probabilistic representation of backtracking counterfactuals is required. However, in our scenario where $\mathbf{X}=\mathbf{F}(\mathbf{U})$ and $\mathbf{F}(.)$ is invertible, a simpler perspective can be adopted. Here, $\mathbf{U}$ can be treated as a deterministic vector, which simplifies the formulation considerably.
@@
Based on this reasoning, we propose the following optimization problem:
@@ 5 Relation Between Backtracking and Interventional Counterfactuals
We consider a SCM $\mathcal{C}=(\mathbf{S},P_{\mathbf{U}})$, where the structural equations $\mathbf{S}$ adhere to an Additive Noise Model (ANM) given by:
$X_{i}:=f_{i}(\mathbf{X}_{\text{pa}(i)})+U_{i},\quad i=1,\dots,n.$ (11)
If the causal model is represented as $\mathbf{X}=\mathbf{F}(\mathbf{U})$, with $\mathbf{F}(.)$ being invertible, the distribution of the noise variables conditioned on $\mathbf{X}=\mathbf{x}$ becomes deterministic. Specifically, all the probability mass of the posterior $\mathbb{P}_{\mathcal{C}}(\mathbf{U}\mid\mathbf{X}=\mathbf{x})$ is assigned to $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$. This determinism is inherent to ANMs due to the invertibility of $\mathbf{F}(.)$.
As outlined in Section 2.2, backtracking counterfactuals aim to produce a desired counterfactual outcome by keeping the causal graph unchanged while minimally modifying the noise variables $\mathbf{U}$ after observing $\mathbf{x}$. Although backtracking counterfactuals are typically defined using the backtracking conditional distribution $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$, the deterministic nature of $\mathbf{U}$ in ANMs eliminates the necessity for statistical modeling. Instead, we directly analyze the connection between backtracking and interventional counterfactuals via their respective causal equations.
@@
In structural causal models adhering to ANM assumptions, backtracking counterfactuals generalize interventional counterfactuals. Specifically, one of the solutions derived from the backtracking counterfactual formulation always coincides with the interventional counterfactual.
@@
Consider a counterfactual query involving a subset of variables $\mathbf{X}^{\mathrm{CF}}_{\mathcal{A}}=\mathbf{x}^{*}_{\mathcal{A}}$. Under the ANM assumption, the posterior distribution $\mathbb{P}_{\mathcal{C}}(\mathbf{U}\mid\mathbf{X}=\mathbf{x})$ assigns probability one to $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$, making $\mathbf{u}$ deterministic. The interventional counterfactuals for this query are defined by the following system of equations:
$\begin{cases}x_{i}^{\mathrm{CF}}=f_{i}(\mathbf{x}_{\text{pa}(i)})+u_{i},&\forall i\notin\mathcal{A},\\ x_{i}^{\mathrm{CF}}=x_{i}^{*},&\forall i\in\mathcal{A}.\end{cases}$ (12)
@@
$\begin{cases}x_{i}^{\mathrm{CF}}=f_{i}(\mathbf{x}_{\text{pa}(i)})+u_{i}^{\mathrm{CF}},&\forall i\notin\mathcal{A},\\ x_{i}^{\mathrm{CF}}=f_{i}(\mathbf{x}_{\text{pa}(i)})+u_{i}^{\mathrm{CF}}=x_{i}^{*},&\forall i\in\mathcal{A}.\end{cases}$ (13)
The key difference between (12) and (13) lies in the adjustment mechanism. Interventional counterfactuals modify the causal graph to enforce $\mathbf{X}^{\mathrm{CF}}_{\mathcal{A}}=\mathbf{x}^{*}_{\mathcal{A}}$, whereas backtracking counterfactuals achieve the same result by adjusting the noise variables. To show that a solution exists for backtracking counterfactuals that aligns with interventional counterfactuals, we define:
$u_{i}^{\mathrm{CF}}=\begin{cases}u_{i},&\forall i\notin\mathcal{A},\\ x_{i}^{*}-x_{i}+u_{i},&\forall i\in\mathcal{A}.\end{cases}$ (14)
Let this modified noise vector be denoted as $\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}}$. Substituting $\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}}$ into the backtracking equations (13) produces equations identical to (12). Thus, interventional counterfactuals are a special case of backtracking counterfactuals.
This reasoning generalizes to any subset of variables $\mathcal{A}$. For any counterfactual query involving $\mathcal{A}$, we can construct $\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}}$ such that backtracking counterfactuals align with interventional counterfactuals. ∎
Theorem 5.1 demonstrates that under the ANM assumption, backtracking counterfactuals inherently include interventional counterfactuals as a specific case. Furthermore, interventional counterfactuals, typically expressed as $\mathbf{x}^{\mathrm{SCF}}=\mathbf{F}_{\mathcal{A}}(\mathbf{u})$, where $\mathbf{F}_{\mathcal{A}}(.)$ represents the structural equations post-intervention, can equivalently be reformulated as $\mathbf{x}^{\mathrm{SCF}}=\mathbf{F}(\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}})$, bridging the gap between the two paradigms.
@@ 6 Connection Between Our Method and Previous Approaches
In our optimization problem (8), setting $\lambda=0$ simplifies the problem to Counterfactual Explanations (4), where causal relationships are disregarded, and the objective becomes finding $\mathbf{x}^{\mathrm{CF}}$ that is closest to $\mathbf{x}$ while modifying the model output.
@@
Assume that the causal model over input variables $\mathbf{x}$ adheres to ANM assumption, the distance functions $d_{X}(\cdot,\mathbf{x})$ and $d_{U}(\cdot,\mathbf{u})$ are convex, $\mathbf{F}(\cdot)$ and $h(\cdot)$ are linear functions, and the cost function is given as $\mathrm{cost}(\mathcal{A};\mathbf{x})=d_{X}(\mathbf{x}^{\mathrm{CF}},\mathbf{x})$. Then, our solution (8) outperforms Causal Algorithmic Recourse. Specifically, for a fixed distance $\alpha$ in the latent space $d_{U}(\mathbf{u}^{\mathrm{CF}},\mathbf{u})=\alpha$, there exists a $\lambda$ such that the solution of (8) yields a counterfactual $\mathbf{x}^{\mathrm{CF}}$ closer to the observed input $\mathbf{x}$ than the solution of (5).
@@
$\begin{split}\arg\min_{\mathbf{x}^{\mathrm{CF}}}\quad&d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)\\ \mathrm{s.t.}\quad&h(\mathbf{x}^{\mathrm{CF}})=y^{\mathrm{CF}},\\ &\mathbf{x}^{\mathrm{CF}}=\mathbf{F}\left(\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}}\right),\\ &\mathbf{x}=\mathbf{F}(\mathbf{u}),\end{split}$ (16)
where $\mathbf{u}_{\mathcal{A}}^{\mathrm{CF}}$ is derived from $\mathbf{u}$ and the intervention set $\mathcal{A}$ (cf. equation (14)).
The main question is whether there exists a $\lambda$ such that the optimal $\mathbf{x}^{\mathrm{CF}}$ in (8) coincides with the optimal $\mathbf{x}^{\mathrm{CF}}$ in (16). Suppose the optimal solution to (16) is attained for $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{CF}}$. Let $\alpha$ represent the distance between $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{CF}}$ and $\mathbf{u}$:
$d_{U}\left(\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{CF}},\mathbf{u}\right)=\alpha.$ (17)
@@
$\begin{split}\arg\min_{\mathbf{x}^{\mathrm{CF}},\mathbf{u}^{\mathrm{CF}}}\quad&d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)\\ \mathrm{s.t.}\quad&h(\mathbf{x}^{\mathrm{CF}})=y^{\mathrm{CF}},\\ &\mathbf{x}^{\mathrm{CF}}=\mathbf{F}\left(\mathbf{u}^{\mathrm{CF}}\right),\\ &d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha,\\ &\mathbf{x}=\mathbf{F}(\mathbf{u}),\end{split}$ (18)
@@
This inequality holds because $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{CF}}$ satisfies the constraint $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$, while other feasible values of $\mathbf{u}^{\mathrm{CF}}$ within the same constraint may reduce the objective $d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)$ further. Hence, the Causal Algorithmic Recourse formulation (16) may not always yield the closest $\mathbf{x}^{\mathrm{CF}}$ to $\mathbf{x}$ among all $\mathbf{u}^{\mathrm{CF}}$ satisfying the distance constraint $\alpha$ from $\mathbf{u}$.
Next, we examine whether there exists a $\lambda$ such that the optimal solution of (8) aligns with the optimal solution of (18). In essence, we seek a $\lambda$ such that the optimal $\mathbf{u}^{\mathrm{CF}}$ from (8) satisfies the distance constraint $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$.
@@
Thus, by varying $\lambda$, it is possible to identify a $\lambda$ such that the solution of (8) matches the solution of (18), ensuring $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$. Consequently, as demonstrated in (19), our proposed method yields a $\mathbf{x}^{\mathrm{CF}}$ that is closer to $\mathbf{x}$ compared to Causal Algorithmic Recourse, while preserving the fixed distance $\alpha$ between the latent variables $\mathbf{u}$ and $\mathbf{u}^{\mathrm{CF}}$. ∎
@@
Importantly, convexity and linearity assumptions are not necessary for the core insight to remain valid. Even without assuming convexity of the distance functions or linearity of $\mathbf{F}(\cdot)$ and $h(\cdot)$—relying solely on the ANM assumption—we can always find a solution that outperforms Causal Algorithmic Recourse among the Pareto optimal points of the vector optimization problem (20). The assumptions of linearity and convexity are only required to guarantee that this Pareto optimal point can be captured by some value of $\lambda$.
Our proposed formulation acts as a framework that bridges and unifies four previously established approaches—Counterfactual Explanations, Deep Backtracking Explanations, Backtracking Counterfactual Explanations, and Causal Algorithmic Recourse. Specifically, when $\lambda=0$, our formulation simplifies to Counterfactual Explanations, which disregard causal relationships. When $\lambda\to\infty$, it corresponds to Deep Backtracking Explanations, focusing exclusively on the latent space. Moreover, our method seamlessly aligns with Backtracking Counterfactual Explanations through the adoption of backtracking conditional distributions. Lastly, under the ANM and convexity assumptions, we demonstrate that our solution outperforms Causal Algorithmic Recourse, generating counterfactuals closer to the original observation while preserving the fixed distance between the latent variables.

### v2

# A New Approach to Backtracking Counterfactual Explanations: A Unified Causal Framework for Efficient Model Interpretability

@@ Abstract
Counterfactual explanations enhance interpretability by identifying alternative inputs that produce different outputs, offering localized insights into model decisions. However, traditional methods often neglect causal relationships, leading to unrealistic examples. While newer approaches integrate causality, they are computationally expensive. To address these challenges, we propose an efficient method called Brace based on backtracking counterfactuals that incorporates causal reasoning to generate actionable explanations. We first examine the limitations of existing methods and then introduce our novel approach and its features. We also explore the relationship between our method and previous techniques, demonstrating that it generalizes them in specific scenarios. Finally, experiments show that our method provides deeper insights into model outputs.
@@ 1 Introduction
Machine learning (ML) has become a core technology in areas such as healthcare, finance, and autonomous systems (Bhoi et al., 2024; Xie et al., 2024; Sancaktar et al., 2022). Although ML models are generally very effective, their limited interpretability is still a significant obstacle (Jethani et al., 2021). Understanding why a model generates a specific prediction is crucial for trust, fairness, and accountability (Miller, 2019; Zhang & Bareinboim, 2018; Von Kügelgen et al., 2022; Karimi et al., 2023). This need is especially clear in high-stakes domains like medical diagnosis or loan approval, where decisions can lead to serious consequences (Doshi-Velez & Kim, 2017).
@@
Despite their benefits, traditional counterfactual methods often overlook causal relationships between features, which can lead to impractical or unrealistic suggestions (Slack et al., 2021). For example, advising someone to lower their income while increasing savings ignores the causal dependency between these factors. This limitation reduces the practical value of such explanations. Causal algorithmic recourse (Karimi et al., 2021) incorporates Interventional Counterfactuals (ICF) to produce more realistic outputs, but this approach is typically computationally expensive and difficult to scale.
Backtracking Counterfactuals (BCF) (Von Kügelgen et al., 2023) present a new way to define counterfactuals in causal inference. We propose a new framework for generating counterfactual explanations using backtracking counterfactuals. Our method combines causal reasoning with computational efficiency, enabling it to produce actionable explanations at scale. The key contributions of our work are as follows:
@@
  • We introduce our novel method, Brace: Backtracking Recourse and Actionable Counterfactual Explanations, that leverages backtracking counterfactuals to provide actionable and meaningful explanations.
@@ 2.1 Structural Causal Models (SCMs)
where $\mathbf{X}_{\text{pa}(i)}$ are the parents variables (direct causes) of $X_{i}$, and $U_{i}$ are independent noise terms sampled from a distribution $P_{\mathbf{U}}$. These relationships are represented by a Directed Acyclic Graph (DAG) $G$, which governs the observational distribution $P_{\mathbf{X}}^{\mathcal{C}}$ (Peters et al., 2017).
@@
commonly referred to as the reduced-form expression. The function $\mathbf{F}(.)$ translates the distribution of latent variables $\mathbf{U}$ into the distribution of observed variables $\mathbf{X}$. We assume causal sufficiency, implying no hidden confounders are present.
Additionally, we adopt a Bijective Generation Mechanism (Nasr-Esfahany et al., 2023), which assumes that $f_{i}(\mathbf{x}_{\text{pa}(i)},\cdot)$ is invertible for fixed $\mathbf{x}_{\text{pa}(i)}$. This ensures the existence of the inverse mapping $\mathbf{F}^{-1}(.)$, allowing us to recover:
@@ 2.2 Interventional and Backtracking Counterfactuals
Both interventional and backtracking perspectives provide valuable insights into counterfactual reasoning but rely on distinct causal reasoning paradigms. Here, we only gave a brief overview of these two approaches. Their precise definitions appear in Appendix A.
@@ 2.3 Problem Definition
The input $\mathbf{x}$ is assumed to follow a SCM $\mathcal{C}=(\mathbf{S},P_{\mathbf{U}})$, where the structural equations $\mathbf{S}$ are fully known. Formally, if $\mathbf{U}$ denotes the latent (noise) variables of the SCM, the input $\mathbf{X}$ is generated as $\mathbf{X}=\mathbf{F}(\mathbf{U})$, with the function $\mathbf{F}(.)$ explicitly defined. The components of $\mathbf{U}$ are mutually independent and $\mathbf{F}(.)$ is invertible. The goal is to find a counterfactual input $\mathbf{x}^{\mathrm{CF}}$ that satisfies:
@@ 4 Our method
In this section, we propose our method called Brace: Backtracking Recourse and Actionable Counterfactual Explanations. As discussed earlier, one of the main limitations of backtracking counterfactuals is their reliance on the conditional distribution $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$. This dependency arises because the choice of $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$ significantly influences the resulting counterfactuals, and its specification is left entirely to the algorithm. Such a distribution is essential when a probabilistic representation of backtracking counterfactuals is required. However, in our scenario where $\mathbf{X}=\mathbf{F}(\mathbf{U})$ and $\mathbf{F}(.)$ is invertible, a simpler perspective can be adopted. Here, $\mathbf{U}$ can be treated as a deterministic vector, which simplifies the formulation considerably.
@@
Based on this reasoning, we propose our method, Brace, with the following optimization problem:
@@ 5 Relation Between Backtracking and Interventional Counterfactuals
Our causal model is represented as $\mathbf{X}=\mathbf{F}(\mathbf{U})$, where $\mathbf{F}(.)$ is an invertible function. Consequently, the distribution of the noise variables conditioned on $\mathbf{X}=\mathbf{x}$ becomes deterministic. Specifically, all the probability mass of the posterior distribution $\mathbb{P}_{\mathcal{C}}(\mathbf{U}\mid\mathbf{X}=\mathbf{x})$ is concentrated at $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$.
As outlined in Section 2.2, backtracking counterfactuals aim to produce a desired counterfactual outcome by keeping the causal graph unchanged while minimally modifying the noise variables $\mathbf{U}$ after observing $\mathbf{x}$. Although backtracking counterfactuals are typically defined using the backtracking conditional distribution $\,\mathbb{P}_{B}(\mathbf{U}^{\mathrm{CF}}\mid\mathbf{U})$, when $\mathbf{F}(.)$ is invertible, the deterministic nature of $\mathbf{U}$ eliminates the necessity for statistical modeling. Instead, we directly analyze the connection between backtracking and interventional counterfactuals via their respective causal equations.
@@
In structural causal models that adhere to the Bijective Generation Mechanism (i.e., $\mathbf{F}(.)$ is invertible), backtracking counterfactuals generalize interventional counterfactuals. Specifically, one of the solutions derived from the backtracking counterfactual formulation always coincides with the interventional counterfactual.
@@
Consider a counterfactual query involving a subset of variables $\mathbf{X}^{\mathrm{CF}}_{\mathcal{A}}=\mathbf{x}^{*}_{\mathcal{A}}$. Under the Bijective Generation Mechanism, the posterior distribution $\mathbb{P}_{\mathcal{C}}(\mathbf{U}\mid\mathbf{X}=\mathbf{x})$ assigns probability one to $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$, making $\mathbf{u}$ deterministic. The interventional counterfactuals for this query are defined by the following system of equations:
$\begin{cases}x_{i}^{\mathrm{ICF}}=f_{i}(\mathbf{x}_{\text{pa}(i)}^{\mathrm{ICF}},u_{i}),&\forall i\notin\mathcal{A},\\ x_{i}^{\mathrm{ICF}}=x_{i}^{*},&\forall i\in\mathcal{A}.\end{cases}$ (11)
@@
$\begin{cases}x_{i}^{\mathrm{BCF}}=f_{i}(\mathbf{x}_{\text{pa}(i)}^{\mathrm{BCF}},u_{i}^{\mathrm{BCF}}),&\forall i\notin\mathcal{A},\\ x_{i}^{\mathrm{BCF}}=f_{i}(\mathbf{x}_{\text{pa}(i)}^{\mathrm{BCF}},u_{i}^{\mathrm{BCF}})=x_{i}^{*},&\forall i\in\mathcal{A}.\end{cases}$ (12)
The key difference between (11) and (12) lies in the adjustment mechanism. Interventional counterfactuals modify the causal graph to enforce $\mathbf{X}^{\mathrm{CF}}_{\mathcal{A}}=\mathbf{x}^{*}_{\mathcal{A}}$, whereas backtracking counterfactuals achieve the same result by adjusting the noise variables.
Due to the DAG assumption in the causal graph, it is clear that equation (11) has a unique solution. After performing interventions on the set $\mathcal{A}$, we obtain multiple DAGs from which we can derive the unique solution for $\mathbf{x}^{\mathrm{ICF}}$ by starting from the source nodes.
Given that $\mathbf{F}(.)$ is invertible, we define
$\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}}=\mathbf{F}^{-1}(\mathbf{x}^{\mathrm{ICF}}).$ (13)
We can also rewrite equation (12) as $\mathbf{x}^{\mathrm{BCF}}=\mathbf{F}(\mathbf{u}^{\mathrm{BCF}})$. By definition, if we substitute $\mathbf{u}^{\mathrm{BCF}}=\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}}$ into the backtracking equations (12), we arrive at the same counterfactual solution, $\mathbf{x}^{\mathrm{ICF}}$. Therefore, interventional counterfactuals can be considered a specific case of backtracking counterfactuals.
This reasoning can be generalized to any subset of variables $\mathcal{A}$. For any counterfactual query involving $\mathcal{A}$, we can construct $\mathbf{u}^{\mathrm{ICF}}$ in such a way that backtracking counterfactuals align with interventional counterfactuals. ∎
To the best of our knowledge, Theorem 5.1 is the first result that relates backtracking and interventional counterfactuals, and it holds independent significance. Theorem 5.1 demonstrates that when $\mathbf{F}(.)$ is invertible, backtracking counterfactuals inherently include interventional counterfactuals as a specific case. Furthermore, interventional counterfactuals, typically expressed as $\mathbf{x}^{\mathrm{ICF}}=\mathbf{F}_{\mathcal{A}}(\mathbf{u})$, where $\mathbf{F}_{\mathcal{A}}(.)$ represents the structural equations post-intervention, can equivalently be reformulated as $\mathbf{x}^{\mathrm{ICF}}=\mathbf{F}(\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}})$, bridging the gap between the two paradigms. By the construction (13) in Theorem 5.1, we can see $\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}}=\mathbf{F}^{-1}\left(\mathbf{F}_{\mathcal{A}}\left(\mathbf{F}^{-1}(\mathbf{x})\right)\right)$.
@@ 6 Connection Between Our Method and Previous Approaches
Our method Brace unifies other existing methods in certain scenarios. In our optimization problem (8), setting $\lambda=0$ simplifies the problem to Counterfactual Explanations (4), where causal relationships are disregarded, and the objective becomes finding $\mathbf{x}^{\mathrm{CF}}$ that is closest to $\mathbf{x}$ while modifying the model output.
@@
Assume that the distance functions $d_{X}(\cdot,\mathbf{x})$ and $d_{U}(\cdot,\mathbf{u})$ are convex, $\mathbf{F}(\cdot)$ and $h(\cdot)$ are linear functions, and the cost function is given as $\mathrm{cost}(\mathcal{A};\mathbf{x})=d_{X}(\mathbf{x}^{\mathrm{CF}},\mathbf{x})$. Then, our method Brace outperforms Causal Algorithmic Recourse. Specifically, for a fixed distance $\alpha$ in the latent space $d_{U}(\mathbf{u}^{\mathrm{CF}},\mathbf{u})=\alpha$, there exists a $\lambda$ such that the solution of (8) yields a counterfactual $\mathbf{x}^{\mathrm{CF}}$ closer to the observed input $\mathbf{x}$ than the solution of (5).
@@
$\begin{split}\arg\min_{\mathbf{x}^{\mathrm{CF}},\ \mathcal{A}}\quad&d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)\\ \mathrm{s.t.}\quad&h(\mathbf{x}^{\mathrm{CF}})=y^{\mathrm{CF}},\\ &\mathbf{x}^{\mathrm{CF}}=\mathbf{F}\left(\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}}\right),\\ &\mathbf{u}_{\mathcal{A}}^{\mathrm{ICF}}=\mathbf{F}^{-1}\left(\mathbf{F}_{\mathcal{A}}\left(\mathbf{u}\right)\right),\\ &\mathbf{x}=\mathbf{F}(\mathbf{u}).\end{split}$ (15)
The main question is whether there exists a $\lambda$ such that the optimal $\mathbf{x}^{\mathrm{CF}}$ in our method (8) coincides with the optimal $\mathbf{x}^{\mathrm{CF}}$ in (15). Suppose the optimal solution to (15) is attained for $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{ICF}}$. Let $\alpha$ represent the distance between $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{ICF}}$ and $\mathbf{u}$:
$d_{U}\left(\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{ICF}},\mathbf{u}\right)=\alpha.$ (16)
@@
$\begin{split}\arg\min_{\mathbf{x}^{\mathrm{CF}},\mathbf{u}^{\mathrm{CF}}}\quad&d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)\\ \mathrm{s.t.}\quad&h(\mathbf{x}^{\mathrm{CF}})=y^{\mathrm{CF}},\\ &\mathbf{x}^{\mathrm{CF}}=\mathbf{F}\left(\mathbf{u}^{\mathrm{CF}}\right),\\ &d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha,\\ &\mathbf{x}=\mathbf{F}(\mathbf{u}).\end{split}$ (17)
@@
This inequality holds because $\mathbf{u}_{\mathcal{A}^{*}}^{\mathrm{ICF}}$ satisfies the constraint $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$, while other feasible values of $\mathbf{u}^{\mathrm{CF}}$ within the same constraint may reduce the objective $d_{X}\left(\mathbf{x}^{\mathrm{CF}},\mathbf{x}\right)$ further. Hence, the Causal Algorithmic Recourse formulation (15) may not always yield the closest $\mathbf{x}^{\mathrm{CF}}$ to $\mathbf{x}$ among all $\mathbf{u}^{\mathrm{CF}}$ satisfying the distance constraint $\alpha$ from $\mathbf{u}$.
Next, we examine whether there exists a $\lambda$ such that the optimal solution of our method (8) aligns with the optimal solution of (17). In essence, we seek a $\lambda$ such that the optimal $\mathbf{u}^{\mathrm{CF}}$ from (8) satisfies the distance constraint $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$.
@@
Thus, by varying $\lambda$, it is possible to identify a $\lambda$ such that the solution of our method Brace (8) matches the solution of (17), ensuring $d_{U}\left(\mathbf{u}^{\mathrm{CF}},\mathbf{u}\right)=\alpha$. Consequently, as demonstrated in (18), our proposed method yields a $\mathbf{x}^{\mathrm{CF}}$ that is closer to $\mathbf{x}$ compared to Causal Algorithmic Recourse, while preserving the fixed distance $\alpha$ between the latent variables $\mathbf{u}$ and $\mathbf{u}^{\mathrm{CF}}$. ∎
@@
Importantly, convexity and linearity assumptions are not necessary for the core insight to remain valid. Even without assuming convexity of the distance functions or linearity of $\mathbf{F}(\cdot)$ and $h(\cdot)$, we can always find a solution that outperforms Causal Algorithmic Recourse among the Pareto optimal points of the vector optimization problem (19). The assumptions of linearity and convexity are only required to guarantee that this Pareto optimal point can be captured by some value of $\lambda$.